OpenStudy (anonymous):

Medal to most helpful in finding correct answer! Using full sentences, describe how to fully simplify Francisco and Ryan's expressions. Describe if Francisco and Ryan started with equivalent expressions or if they started with expressions that are not equal.

OpenStudy (anonymous):

Francisco and Ryan are stuck simplifying radical expressions. Francisco has simplified to \[\frac{ x ^{1/2} }{ x ^{3/18} }\]. Ryan has simplified to \[\sqrt[27]{x^{2}\times x^{3}\times x^{4}}\]. Using full sentences, describe how to fully simplify Francisco and Ryan's expressions. Describe if Francisco and Ryan started with equivalent expressions or if they started with expressions that are not equal.

OpenStudy (dariusx):